A square garden is to be enlarged to a rectangular by adding 5 ft. to the length and deducting 4 ft. from the width, with area unchanged. Determine present and new dimensions.

My partial attempt:
Area = length * width
Let s = side of square.
Square:
Length: s Width: s Area: s^2
Rectangle:
Length: s + 12 Width: s - 4 Area: (s + 5) (s - 4)

Not sure how to proceed. Thanks.

Answer by ewatrrr(24785)

Area = length * width
Let s = side of square. A(square) = s^2
Square:
Length: s Width: s Area: s^2
Rectangle:
Length: s + 12 Width: s - 4
Area: (s + 5) (s - 4)
area unchanged.
(s + 5) (s - 4) = s^2
s^2 + s - 20 = s^2
s = 20ft for the Square
25ft by 16ft the Rectangle
AND
400ft^2 = 25ft*16ft CHECKS!